# End-to-end Streaming model for Low-Latency Speech Anonymization

### v1

@@ 5.4 Speaker Identity transfer
In a final step, we evaluated our models’ ability to capture the voice of a target speaker. For this purpose, we used an objective score of speaker similarity based on the cosine similarity between speaker embeddings of the target and the synthesized utterances [^7] Computed using: https://github.com/resemble-ai/Resemblyzer. We compare our model against the three VC baselines (VQMIVC [11], QuickVC [37], DiffVC [38]) using the same settings as those in section 5.2) to generate VC samples. Results are summarized in the rightmost column of Table 2 (SSS). As a guideline, pairing two utterances from the same speaker yields an average cosine similarity of 0.89. As shown, our base model outperforms the three baselines, achieving cosine similarity that it close to the average within-speaker similarity of 0.89.
@@ 6 Discussion
Most existing speaker anonymization methods do not operate in low-latency streaming mode, preventing their use in field operations. In this paper, we present an end-to-end streaming model that operates with low latency and achieves anonymization by mapping speaker embedding into an artificially generated pseudo speaker in a causal fashion (i.e., no future context). While there exists a quality-latency tradeoff, our system can achieve latency s low as 66ms while maintaining state-of-the-art naturalness, intelligibility and privacy preservation. Our lite version is roughly 10MB and can potentially be deployed on mobile devices to support real-time field applications. Accent can carry speaker related cues [41] and in future work, we aim to add accent conversion to this pipeline. Other research direction is to add the control of emotion while synthesizing speech signals.

### v2

@@ 5.4 Speaker Identity transfer
In a final step, we evaluated our models’ ability to capture the voice of a target speaker. For this purpose, we used an objective score of speaker similarity based on the cosine similarity between speaker embeddings of the target and the synthesized utterances obtained from a ASV system [^7] Computed using: https://github.com/resemble-ai/Resemblyzer. We compare our model against the three VC baselines (VQMIVC [11], QuickVC [39], DiffVC [40]) using the same settings as those in section 5.2) to generate VC samples. Results are summarized in the rightmost column of Table 2 (SSS). As a guideline, pairing two utterances from the same speaker yields an average cosine similarity of 0.89. As shown, our base model outperforms the three baselines, achieving cosine similarity that is close to the average within-speaker similarity of 0.89.
@@ 6 Discussion
Most existing speaker anonymization methods do not operate in low-latency streaming mode, preventing their use in field operations. In this paper, we present an end-to-end streaming model that operates with low latency and achieves anonymization by mapping speaker embedding into an artificially generated pseudo speaker in a causal fashion (i.e., no future context). The pseudo-speaker generator can produce speaker embeddings that are very close to a real person in the corpus. Although we train the pseudo-speaker generator on a different corpus than the speech anonymization system to guard against this possibility, we could also test if a pseudo-speaker is too close to one on the corpus or outside the space of speakers in the training corpus, and generate new ones until a valid one is generated. While there exists a quality-latency tradeoff, our system can achieve latency as low as $66ms$ while maintaining state-of-the-art naturalness, intelligibility and privacy preservation. Our lite version is roughly 10MB and can potentially be deployed on mobile devices to support real-time field applications. Accent can carry speaker related cues [43] and in future work, we aim to add accent conversion to this pipeline. Other research direction is to add the control of emotion while synthesizing speech signals.